# Graph-Enhanced EEG Foundation Model

### v1

@@ Results
Table 3: The results for all downstream tasks and the GNN architectures we experimented with using padding as the sequence length adjuster. GCN and GAT without edge weights were excluded from these experiments due to their poor performance with the linear layer (see Table 2). Entries from our proposed models that exceed the baseline are underlined.
Comparison of Base Model Configurations (RQ4)
This section examines whether incorporating GNNs is more effective when the base model configuration is “BENDR” or “Linear.” As shown in Tables 2 and 3, when comparing the performance of models employing GNNs across all task-model combinations, the “BENDR” configuration outperformed the “Linear” configuration in 21 out of 24 cases. In contrast, among the baseline models, “Linear” exceeded “BENDR” in 2 out of 3 cases. Notably, all GNN-based models that outperformed the baseline were configured with “BENDR.” These results indicate that incorporating GNNs is more effective when the base model configuration is “BENDR.”
@@
Conclusion
Foundation models for EEG analysis are particularly valuable due to the difficulty of collecting large amounts of labeled data and their ability to be applied to a wide range of EEG tasks with minimal computational time and cost. In this study, we propose an improved foundation model for EEG that leverages both inter-channel relationships and the temporal dynamics of EEG signals. The proposed architecture integrates GNNs, which are effective at learning relationships between entities, with a masked autoencoder-based framework. We evaluated the model using several GNN architectures across three downstream tasks. The results indicate that our proposed model, when employing the GCN architecture (Kipf and Welling 2016) with specific configurations, consistently outperformed the baseline across all tasks. These findings demonstrate that incorporating inter-channel relationships learning through GNNs enhances the model’s performance, establishing it as a more effective foundation model for EEG analysis.
As future work, we plan to evaluate the proposed model on a broader range of tasks to further demonstrate its versatility. Additionally, we aim to investigate the mechanisms underlying the observed improvements achieved through integrating graph structures, using techniques such as GNNExplainer (Ying et al. 2019). And another potential direction is to expand our approach, which integrates inter-channel relationships and time-series information, to other base models or various self-supervised learning approaches Alternatively, we may design a novel architecture specifically optimized for EEG foundation models.
Acknowledgments

### v2

@@ IV-C3 Comparison of Sequence Length Adjusters (RQ3)
TABLE III: The results for all downstream tasks and the GNN architectures we experimented with using padding as the sequence length adjuster.
IV-C4 Comparison of Base Model Configurations (RQ4)
This section examines whether incorporating GNNs is more effective when the base model configuration is “BENDR” or “Linear.” As shown in Tables II and III, when comparing the performance of models employing GNNs across all task-model combinations, the “BENDR” configuration outperformed the “Linear” configuration in 21 out of 24 cases. In contrast, among the baseline models, “Linear” exceeded “BENDR” in 2 out of 3 cases. Notably, all GNN-based models that outperformed the baseline were configured with “BENDR.” These results indicate that incorporating GNNs is more effective when the base model configuration is “BENDR.”
@@ IV-C4 Comparison of Base Model Configurations (RQ4)
V Conclusion
Foundation models for EEG analysis are particularly valuable due to the difficulty of collecting large amounts of labeled data and their ability to be applied to a wide range of EEG tasks with minimal computational time and cost. In this study, we propose an improved foundation model for EEG that leverages both inter-channel relationships and the temporal dynamics of EEG signals. The proposed architecture integrates GNNs, which are effective at learning relationships between entities, with a masked autoencoder-based framework. We evaluated the model using several GNN architectures across three downstream tasks. The results indicate that our proposed model, when employing the GCN architecture [7] with specific configurations, consistently outperformed the baseline across all tasks. These findings demonstrate that incorporating inter-channel relationships learning through GNNs enhances the model’s performance, establishing it as a more effective foundation model for EEG analysis.
As future work, we plan to evaluate the proposed model on a broader range of tasks to further demonstrate its versatility. Additionally, we aim to investigate the mechanisms underlying the observed improvements achieved through integrating graph structures, using techniques such as GNNExplainer [26]. And another potential direction is to expand our approach, which integrates inter-channel relationships and time-series information, to other base models or various self-supervised learning approaches Alternatively, we may design a novel architecture specifically optimized for EEG foundation models.
VI Acknowledgments